OpenStudy (anonymous):

I will give medals to all and thank you personally on your page (comment wall) check this word problem out. it's two parts. goodluckk!!

OpenStudy (anonymous):

You are familiar with the following types of factoring:•factoring out the Greatest Common Factor (GCF). •factoring by grouping . •factoring trinomials of the form x^2 + bx + c and ax^2 + bx + c. As you know, you need to know the first two types of factoring listed above in order to be successful in factoring trinomials of the form ax^2 + bx + c. Part 1: In your own words, explain how a trinomial of the form 2x^2 + 13x + 15 can be turned into a four term polynomial suitable for factoring by grouping. Use complete sentences. Part 2: If you were an Algebra 1 instructor and were creating a test on factoring trinomials of the form ax^2 + bx + c, what do you think would be the easiest way to create a trinomial that can be factored? Provide one unique example.

OpenStudy (anonymous):

part 2. start in factored form say as \[(2x-1)(x+3)\] then multiply out

OpenStudy (anonymous):

so part 2 is 2 x^2 + 5 x-3

OpenStudy (anonymous):

thats is my example yes. you can make up any one you like

OpenStudy (anonymous):

ok so part two would be (2x-1) (x+3) equals out to be 2x^2+5x-3

OpenStudy (anonymous):

part 1 ?

OpenStudy (anonymous):

\[2x^2+13x+15=2x^2+10x+3x+15=2x(x+5)+3(x+5)=(2x+3)(x+5)\] the words are on you

OpenStudy (anonymous):

:/ THATS A BIT DIFFICULT.

OpenStudy (anonymous):

hero ? a little help